Jack is 2 times as old as Lacy. 8 years from now the sum of their ages will be 55. How old are they now?
Answer by Maths68(1474)


Let
Present age of Jack = j
Present age of Lacy = l

Jack is 2 times as old as Lacy.
j=2l........(1)
8 years from now
Jack age will be = j + 8
Lacy age will be = l +8
Then, the sum of their ages will be 55
j+8+l+8=55
j+l+16=55
j+l=55-16
j+l=39............(2)
Put the value of j from (1) to (2)
2l+l=39
3l=39
l=39/3
l=13
Put the value of l in (1)
j=2l
j=2(13)
j=26



Present age of Jack = j = 26
Present age of Lacy = l = 13